OpenStudy (anonymous):

a can of beans has surface area 359 cm^2. Its height is 10 cm. What is the radius of the circular top?

OpenStudy (anonymous):

Do you know the formula for calculating the surface area of a sphere?

OpenStudy (anonymous):

2pir^2+2pirh?

OpenStudy (anonymous):

oh wait just pir^2?

OpenStudy (anonymous):

Pir^2 is just the surface area of one lid, so for the 2 lids it would be 2Pir^2 what is the formula for the walls?

OpenStudy (anonymous):

2pirh?

OpenStudy (anonymous):

is the answer 4.06?

OpenStudy (anonymous):

im following the steps the teacher gave me for notes but i keep getting them wrong. This is the 5th problem like this I've attempted to do.

OpenStudy (anonymous):

The wall of the can would be circumference x height. (The height is given). What is the circumference?

OpenStudy (anonymous):

I did not calculate the answer. I'm doing step by step with you.

OpenStudy (anonymous):

I apologize, you wrote the correct answer earlier. The surface area of the wall is 2 pirh.

OpenStudy (anonymous):

And what you wrote the very first time "2pir^2+2pirh" is also correct.

OpenStudy (anonymous):

359=2pir^2 + 20pir?

OpenStudy (anonymous):

So far looks good. It may be easier to factor out a 2 and write it as 359 = 2 (pir^2 + 10pir). You're on the right track.

OpenStudy (anonymous):

in my notes my teacher has me equal it out to 0 following this it would be 0 = 2pir^2+20pir-359

OpenStudy (phi):

this one is more painful than that. How about divide all term by 2pi and write it as \[ r^2 + 10 r - \frac{359}{2 \pi} = 0 \] if we use a calculator, that is about \[ r^2 + 10 r - 57.1366=0 \] to find r we could use the quadratic formula

OpenStudy (anonymous):

after that step i'm getting very confused

OpenStudy (anonymous):

Give me a few minutes. I'll try it the way you have it in your notes.

OpenStudy (anonymous):

x=-b+- sqr rt, b^2-4ac / 2a

OpenStudy (anonymous):

quadratic formula?

OpenStudy (anonymous):

okay thank you

OpenStudy (phi):

yes, quadratic formula. if you first divide by 2pi most of the numbers get a bit smaller and you would have a=1, b=10 , c= 57.1366

OpenStudy (anonymous):

-10pi+- sq rt. 100-4(!)(57.1366)/2?

OpenStudy (phi):

if you are doing \[r^2 + 10 r - 57.1366=0 \] -b is -10 (not -10pi) a=1 so 4ac is just 4*-57.1366 (notice c is negative)

OpenStudy (phi):

so \[ \frac{-10+ \sqrt{100 + 228.5465}}{2} \]

OpenStudy (phi):

I would pick the + sqrt because otherwise we get a negative radius

OpenStudy (anonymous):

114.27325

OpenStudy (anonymous):

am i calculating this right

OpenStudy (anonymous):

i feel so wrong

OpenStudy (phi):

yes, not correct.

OpenStudy (anonymous):

what am i doing wronnggggg

OpenStudy (phi):

can you get this far: \[ \frac{-10+ \sqrt{100 + 228.5465}}{2}\] (by the way for c = -359/(2pi) I would use all the digits the calculator has)

OpenStudy (anonymous):

-28.56831228

OpenStudy (phi):

for \[ r^2 + 10 r - 57.1366=0 \] a=1, b=10, c= -57.1366 (roughly) let's first figure out b^2 -4 * a *c that is 100 - 4 * 1 * -57.1366 what does that simplify to ?

OpenStudy (anonymous):

i'm getting -118.5465.

OpenStudy (phi):

4*-57.1366= -228.5465 now do 100 - (-228.5465)

OpenStudy (anonymous):

I made a mistake in my calculation. I don't want to post until I figure out where the error is.

OpenStudy (phi):

notice it is minus a minus. that makes it a +

OpenStudy (anonymous):

OHHHHHHH, 328.5465

OpenStudy (anonymous):

i forgot the other minus

OpenStudy (phi):

yes. so make a note: Be careful about the minus signs (they are evil) now find the square root.

OpenStudy (anonymous):

4.062925852??????

OpenStudy (phi):

yes, that looks good. do they want the answer rounded ?

OpenStudy (anonymous):

yes, 4.06

OpenStudy (anonymous):

so i was right earlier. Thank you both so much for your help!

OpenStudy (anonymous):

Sorry I wasn't able to help you through it

OpenStudy (phi):

if we use 4.06 in the original formula we get 358.67 or about 359, so it checks out.

OpenStudy (anonymous):

If it helps I'll post how to do it in a calculator that has fractions

OpenStudy (anonymous):

\[\frac{ -20 + \sqrt{20^2+ 8(\frac{ 359 }{ \pi })} }{ 4 }\]

OpenStudy (anonymous):

If your calculator allows you to enter the above equation (in one step) you'll get the same answer. I put in + 8 since the two negatives become positive.